Times-Roman Publishing Company reports the following amounts in his first three years of operation:
($ in 000s)           2018.    2019.     2020
pretax acct inc.   $200.     $180.     $170
taxable income.    220.        190.       180

the difference between pre-tax accounting income and taxable income is due to subscription revenue for one year magazine subscriptions being reported for tax purposes in the year received, but reported in the income statement in later years when the performance obligation is satisfied. The income tax rate is 40% each year. Times- Roman anticipates profitable operations in the future.

1. what is the balance sheet account for Wichita temporary difference is created by this situation?

2. for each year, indicate the cumulative amount of the temporary difference at year end.

3. determine the balance in the related deferred tax account at the end of each year. Is it a deferred tax asset or deferred tax liability?
Answer #1
Solution 1:
The balance sheet account is "Unearned subscription revenue" for which a temporary difference is created by this situation.The unearned subscription is created when the services are not performed but the advance payment is created by service providing party.
Solution 2:
Times Roman Publishing company
Computation of Cumulative Temporary differences at year end (In '000)
Year Pre tax accounting income Taxable Income Deductible/(Taxable) temporary differences for the year Cumulative temporary differences at year end
2018 $200.00 $220.00 $20.00 $20.00
2019 $180.00 $190.00 $10.00 $30.00
2020 $170.00 $180.00 $10.00 $40.00
Solution 3:
As these are deductible temporary differences, therefore these difference give rise to deferred tax assets:
Balance in deferred tax account at the end of each year:
2018 = $20 * 40% = $8
2019 = $30 * 40% = $12
2020 = $40*40% = $16